• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 16:43:28 on September 26, 2005

    Entry number 156042

    keyword=Reduced Helium target density by 23%

    I have reduced the density of the 4He target by bleeding off gas. We are now at 106 psi and T = 6.70 K, which corresponds to a density of about 0.103 g/cc (compared to our standard operating point for this run of 0.134 g/cc), so about a 23% reduction.

    Once the septa are fixed (the scanner work is done), we can do spectrometer DAQ runs to look at the Al window background; we can also run the scanner to see if the infamous "shelf" changes in any way.

    Kirsten and Lisa are out of the hall, having re-installed the scanner pre-radiators.